OpenStudy (anonymous):

solve the system by using a table. -0.2x-0.3y=0.2 0.4x-0.3y=1.4 answers: (2,2) (2,-2) (2,1) (-2,2)

OpenStudy (anonymous):

Can u explain what u mean by table

OpenStudy (anonymous):

you have to use a graphing calculator to solve

OpenStudy (anonymous):

U can try this method i forget what it is called Setting X or Ys to equal the same digit. For X Multiplying first equation by 2 so you have -0.4x-0.6y=0.4 now add down so its looks like this \[-0.4x-0.6y=0.4 \] \[+\] \[-0.4x-0.3y=1.4 \] So equation is \[-0.9y=1.8\] Now solve for y \[y=-2\] Now plug back in either equation \[-0.4x-0.6*(-2)=0.4 \] = \[-0.4x+1.2=0.4 \] Now solve for x by first subtracting 1.2 then dividing by 0.4 \[-0.4x=0.4-1.2\] \[\frac{ -0.4x }{ 0.4 } x=\frac{ -0.8 }{ -0.4 }\] \[x=2\] Therefore , we conclude it to be \[(2,-2)\]

OpenStudy (anonymous):

thank you so much!!